Determine the coordinates of the x-intercepts for the quadratic equation 4x2 – 3x - 1 = 0.
Answer by ewatrrr(24785)


Hi
finding the coordinates of the x-intercepts
4x^2 – 3x - 1 = 0
factoring
(4x + 1)(x-1)= 0
(4x + 1)=0
x = -(1/4)
(x-1)= 0
x = 1
coordinates of the x-intercepts are(1,0) and (-.25,0)